# NVLM: Open Frontier-Class Multimodal LLMs

### v1

@@ Abstract
To advance research in the field, we are releasing the model weights and will open-source the code for the community: https://nvlm-project.github.io/.
@@ 1 Introduction
Figure 2: NVLM 1.0 versus leading proprietary and open-access multimodal LLMs. Note that the model weights for Llama 3-V have not been released as of the time of this report. The results demonstrate that NVLM 1.0 achieves performance on par with leading models across both vision-language and text-only tasks. Additionally, we compare multimodal LLM to its backbone LLM on text-only tasks. Llama 3-V 70B and 405B show no degradation in text-only tasks, as their LLM backbones are frozen during multimodal training. In contrast, our NVLM-D1.0 72B model demonstrates significant improvements over its text backbone on text-only math and coding benchmarks, with average accuracy increasing by 4.3 points after multimodal training. See Table 7 and Table 8 for full details.
@@ 2. $\bullet$
    Model designs that handle high-resolution image input (e.g, dynamic high-resolution [80; 30; 18]) significantly boost performance on OCR-related tasks (e.g., OCRBench [81]), but sometimes show reduced accuracy on reasoning-related tasks (e.g., MMMU [166]) compared to their low-resolution counterparts.
@@ 2 Qualitative Study
Our NVLM-D1.072B demonstrates versatile capabilities in various multimodal tasks by jointly utilizing OCR, reasoning, localization, common sense, world knowledge, and coding ability. For instance, our model can understand the humor behind the “abstract vs. paper” meme in Figure 1 (a) by performing OCR to recognize the text labels for each image and using reasoning to grasp why juxtaposing “the abstract” — labeled with a fierce-looking lynx — and “the paper” — labeled with a domestic cat — is humorous. NVLM accurately performs localization to effectively answer location-sensitive questions, such as “What is the difference between the left, middle, and right objects in the image?” in Figure 1 (b). NVLM is capable of performing mathematical reasoning and coding based on visual information, such as tables and handwritten pseudocode, as illustrated in Figure 1 (d) and (e). For more examples, refer to Appendix A.
@@ 3.1 Essential Building Blocks
Multimodal LLMs [e.g., 4; 65; 67; 19] typically leverage pretrained vision encoders (e.g., CLIP [120]) to extract visual features from input images or video frames, with only a very few exceptions [3]. These vision encoders [120; 47; 19; 167; 27] are often trained on large-scale, diverse, and noisy text-image pairs sourced from the web [124; 12; 33]. This allows for large-scale training and enhances the generalization needed to effectively process visual input in unseen domains. The other types of datasets, such as those used for optical character recognition (OCR) [19] and image segmentation [58], are also incorporated to enhance the specific capabilities of vision encoders. In this study, we use InternViT-6B [19] as the default vision encoder due to its strong performance. We keep this vision encoder frozen at all stages of training, as this simplifies the training process while still delivering strong results.
@@ 3.3 High-Resolution Inputs
Properly handling high-resolution images is crucial for achieving strong performance in many OCR-related tasks. However, vision encoders are typically trained with static resolution of $224^{2}$ or $336^{2}$ pixels for efficiency [120; 100], when the image patch size per token is usually $14^{2}$ or $16^{2}$. For example, feeding a $224^{2}$ image to ViT-L/14 (patch size $14^{2}$) results in $(\frac{224}{14})^{2}=256$ tokens. There are specialized vision encoders that can directly handle static high-resolution images. For instance, the SAM encoder [58], designed for image segmentation, can process images of $1024^{2}$ pixels with a ViT-L/16 backbone ($16^{2}$ pixels per patch), producing a 4096-token output. This can be costly, especially when training datasets and downstream tasks contain a mix of low-resolution and high-resolution images.
@@ 4.1 Shared Vision Pathway
This dynamic high-resolution (DHR) design significantly improves performance on OCR-related tasks [18; 30], but sometimes results in degraded performance on reasoning-related tasks [166] when all image tokens from the tiles are simply concatenated and fed directly into the LLM. We will address this issue across the three architectures, respectively.
@@ 5.2 Multimodal SFT Data
Following previous leading open-source work [e.g., 18; 65], we incorporate the training splits of datasets including ChartQA [93], DocVQA [95], VQAv2 [38], TextVQA [131] and AI2D [55] into the SFT blend. Their test sets are used as evaluation benchmarks in Section 6, meaning they are not evaluated in a zero-shot setting. Note that it is unknown whether the proprietary multimodal LLMs are being evaluated on these benchmarks in a zero-shot or fine-tuning setting, as no information is provided regarding their training datasets.
@@ 7 Conclusion
We introduce NVLM-1.0, a family of frontier multimodal large language models that achieve state-of-the-art results on vision-language tasks, rivaling leading multimodal LLMs, without compromising text-only performance during multimodal training. Furthermore, we provide key insights on architecture design, tile-based dynamic high-resolution input, multimodal training data curation, and how to achieve production-grade multimodality with even improved text-only performance after multimodal training. We open-source the model weights and provide the full technical details to the community.

### v2

@@ Abstract
To advance research in the field, we release the model weights at https://huggingface.co/nvidia/NVLM-D-72B and will open-source the training code for the community soon. For more information, please visit our project page at https://research.nvidia.com/labs/adlr/NVLM-1.
@@ 1 Introduction
Figure 2: NVLM 1.0 versus leading proprietary and open-access multimodal LLMs. Note that the model weights for Llama 3-V [82] have not been released as of the time of this report. The results demonstrate that NVLM 1.0 achieves performance on par with leading models across both vision-language and text-only tasks. Additionally, we compare multimodal LLM to its backbone LLM on text-only tasks. Llama 3-V 70B and 405B show no degradation in text-only tasks, as their LLM backbones are frozen during multimodal training. In contrast, our NVLM-D1.0 72B model demonstrates significant improvements over its text backbone on text-only math and coding benchmarks, with average accuracy increasing by 4.3 points after multimodal training. See Table 7 and Table 8 for full details.
@@ 2. $\bullet$
    Model designs that handle high-resolution image input (e.g, dynamic high-resolution [160; 80; 30; 18]) significantly boost performance on OCR-related tasks (e.g., OCRBench [81]), but sometimes show reduced accuracy on reasoning-related tasks (e.g., MMMU [166]) compared to their low-resolution counterparts.
@@ 2 Qualitative Study
Our NVLM-D1.072B demonstrates versatile capabilities in various multimodal tasks by jointly utilizing OCR, reasoning, localization, common sense, world knowledge, and coding ability. For instance, our model can understand the humor behind the “abstract vs. paper” meme in Figure 1 (a) by performing OCR to recognize the text labels for each image and using reasoning to grasp why juxtaposing “the abstract” — labeled with a fierce-looking lynx — and “the paper” — labeled with a domestic cat — is humorous. NVLM accurately performs localization to effectively answer location-sensitive questions, such as “What is the difference between the left, middle, and right objects in the image?” in Figure 1 (b). NVLM is capable of performing mathematical reasoning and coding based on visual information, such as tables and handwritten pseudocode, as illustrated in Figure 1 (d) and (e). For more examples, refer to Appendix A or our project site: https://nvlm-project.github.io/.
@@ 3.1 Essential Building Blocks
Multimodal LLMs [e.g., 4; 65; 67; 19] typically leverage pretrained vision encoders (e.g., CLIP [120]) to extract visual features from input images or video frames, with only a very few exceptions [3]. These vision encoders [120; 47; 19; 167; 27] are often trained on large-scale, diverse, and noisy text-image pairs sourced from the web [124; 12; 33]. This allows for large-scale training and enhances the generalization needed to effectively process visual input in unseen domains. The other types of datasets, such as those used for optical character recognition (OCR) [19] and image segmentation [58], are also incorporated to enhance the specific capabilities of vision encoders. In this study, we use InternViT-6B [19] as the default vision encoder due to its strength. We keep this vision encoder frozen at all stages of training, as this simplifies the training process while still delivering strong results.
@@ 3.3 High-Resolution Inputs
Properly handling high-resolution images is crucial for achieving state-of-the-art results in many OCR-related tasks. However, vision encoders are typically trained with static resolution of $224^{2}$ or $336^{2}$ pixels for efficiency [120; 100], when the image patch size per token is usually $14^{2}$ or $16^{2}$. For example, feeding a $224^{2}$ image to ViT-L/14 (patch size $14^{2}$) results in $(\frac{224}{14})^{2}=256$ tokens. There are specialized vision encoders that can directly handle static high-resolution images. For instance, the SAM encoder [58], designed for image segmentation, can process images of $1024^{2}$ pixels with a ViT-L/16 backbone ($16^{2}$ pixels per patch), producing a 4096-token output. This can be costly, especially when training datasets and downstream tasks contain a mix of low-resolution and high-resolution images.
@@ 4.1 Shared Vision Pathway
This dynamic high-resolution (DHR) design significantly improves performance on OCR-related tasks [18; 30], but sometimes results in degraded results on reasoning-related tasks [166] when all image tokens from the tiles are simply concatenated and fed directly into the LLM. We will address this issue across the three architectures, respectively.
@@ 5.2 Multimodal SFT Data
Following previous leading open-source work [e.g., 18; 65], we incorporate the training splits of datasets including ChartQA [93], DocVQA [95], VQAv2 [38], TextVQA [131] and AI2D [55] into the SFT blend. Their test sets are used as evaluation benchmarks in Section 6, meaning they are not evaluated in a zero-shot setting. We want to emphasize that we did not apply data augmentation to the training splits of these benchmark datasets. While such techniques certainly improve benchmark results, they contribute minimal value to the model’s out-of-domain generalization. Note that it is unknown whether the proprietary multimodal LLMs are being evaluated on these benchmarks in a zero-shot or fine-tuning setting, as no information is provided regarding their training datasets. We hypothesize that it is a fine-tuning setting, based on observed accuracy gaps between the training and test sets for some proprietary models; however, this is not conclusive.
@@ 7 Conclusion
We introduce NVLM-1.0, a family of frontier multimodal large language models that achieve state-of-the-art results on vision-language tasks, rivaling leading multimodal LLMs, without compromising text-only performance during multimodal training. Furthermore, we provide key insights on architecture design, tile-based dynamic high-resolution input, multimodal training data curation, and how to achieve production-grade multimodality with even improved text-only performance after multimodal training. We open-source the model weights and training code to the community. Please refer to the updated information at https://nvlm-project.github.io/.